OpenStudy (anonymous):

If t is a prime number, and x is a factor of 12, then t/x could be all of the following EXCEPT? A) 1/12 B) 1/4 C) 1/2 D) 1 E) 2

OpenStudy (anonymous):

factor 12....

OpenStudy (anonymous):

i think the point is that if t is prime, it is not divisible by 2 or 3 or 6 or 12

OpenStudy (anonymous):

so \[\frac{t}{x}\] cannot be 2

OpenStudy (anonymous):

I'm not sure how it can be 1/12 either...unless they are counting 1 as a prime (which it's not)

OpenStudy (anonymous):

maybe i am wrong though.

OpenStudy (anonymous):

ok it certainly could be 1 yes?

OpenStudy (anonymous):

I agree with the answer of 2, satellite :)

OpenStudy (anonymous):

t=3, x=3 gives you 1 t=2, x=4 gives you 1/2 t=3, x=12 gives you 1/4 That's all I can "prove"

OpenStudy (anonymous):

1 is not a prime...

OpenStudy (anonymous):

no i meant the ratio could be 1

OpenStudy (anonymous):

Ahh... t=2, x=1 gives you 2!!!

OpenStudy (anonymous):

Answer is 1/12

OpenStudy (anonymous):

silly me..not thinking about 1 as a factor of 12

OpenStudy (anonymous):

yeah well silly me too

OpenStudy (anonymous):

\[\frac{t}{x}=\frac{1}{12}\] \[12t=x\] not likely

OpenStudy (anonymous):

Well, since t > 1, in order for t/x = 1/12, it must be that x > 12. But since x is a factor of 12, x <= 12. Which is a contradiction. :)